OpenStudy (anonymous):

Solve 1/v = 9/v - 1/5

jimthompson5910 (jim_thompson5910):

1/v = 9/v - 1/5 5 = 45 - v ... Multiply EVERY term by the LCD 5v to clear out the fractions 5-45 = -v -40 = -v -40/(-1) = v 40 = v So v = 40